### v2

# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation via Large-Scale Pseudo Labelling

                             Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth to compare and filter bad examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distil models in low-resource settings. To address this challenge, we propose an unsupervised or label-free framework for distillation, thus eliminating the requirement for labeled data altogether. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER. Additionally, our models are on par with or better than similar supervised data filtering setup. When we scale the data, our models significantly outperform all zero-shot and supervised models. We demonstrate that it is possible to distill large Whisper models into relatively small models without using any labeled data. Our distilled models are 25-50% more compute and memory efficient while maintaining performance equal to or better than the teacher model.
1 Introduction
In an effort to democratize Automatic Speech Recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023b); Radford et al. (2023); Communication et al. (2023a); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to counterparts like English. For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows a performance in the single digits on English and Spanish, East Asian and African languages show a much lower performance [^1] https://github.com/openai/whisper/discussions/1762. While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-hungry and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2019); Frantar et al. (2022). It has proven to be powerful, effectively reducing model size (and hence compute and memory requirements) while keeping performance on par with the model originally distilled from. Gandhi et al. (2023a) and Waheed et al. (2024) showcase this for English and Arabic, respectively. That being said, one important limitation of these works is their reliance on a large number of high-quality pseudo labels from the teacher model. Ground-truth labels, which are scarce for low-resource languages, are needed to filter out incorrect teacher predictions and ensure such quality. This constraint thus calls for an unsupervised approach to knowledge distillation.
In this work, we propose frameworks to distill knowledge from teacher models that do not rely on filtered pseudo-labels. Our method achieves a similar performance to supervised distillation and outperforms pseudo-label distillation without filtering. Our contributions are the following:
  • We explore various unsupervised methods to filter-out low-quality pseudo-labels, eliminating the necessity of labeled data for distillation.
  • We assess the performance of our distilled models across diverse settings and extend the existing evaluation to include the SADA dataset.
  • We analyze the effectiveness of these metrics in detecting low-quality pseudo-labeled examples.
We organize the paper as follows: we present a review of the existing literature in Section 2 before introducing knowledge distillation along with our methodology in Section 3. Section 4 contains particulars of our experiments and Section 5 delves into the results we obtained. We discuss these results in Section 5, and go over our limitations and conclude our work in Sections 7 and 6, respectively.
2 Related Work
Knowledge distillation is the process through which knowledge from a trained large model referred to as the teacher is transferred to a smaller model called the student, in order to minimize the requirements at both the compute and memory levels Gandhi et al. (2023a); Gou et al. (2021); Lopes et al. (2017a). Different approaches for knowledge distillation have been explored such as Patient Knowledge Distillation (PKD) Sun et al. (2019), T-S learning Manohar et al. (2018), Cross-Modal Hashing (CMH) Hu et al. (2020) joint Unsupervised Domain Adaptation with KD Zhang et al. (2021) and Data-Free KD Lopes et al. (2017b) among others in the NLP and CV fields alike. Various works have showcased its use specifically for speech recognition, such as Shao et al. (2023) who shrink a Whisper Radford et al. (2023) model tangibly (around 80% of its original size), all while improving its performance, and a Hubert model is reduced by Chang et al. (2021) to 75% of its original size through layer-wise distillation without a significant drop in performance.
Thanks to the rise of multilingual models such as Whisper Radford et al. (2023), Massively Multilingual Speech (MMS) Pratap et al. (2023b), and Seamless-M4T Communication et al. (2023b), low-resource ASR is garnering more attention. Additionally, pseudo-label knowledge distillation is showing promise in the low-resource domain by leveraging these multilingual models to create tailored students that are superior in performance. That being said, the issue of data scarcity becomes more pressing with the reliance of this distillation approach on a large number of ground-truth transcriptions, effectively creating a bottleneck in the pipeline. Doutre et al. (2021) showcase an approach to unsupervised knowledge distillation where they use non-streaming models to distill knowledge into streaming models (which are known to perform worse). Inspired by Liao et al. (2013), they use randomly-segmented audio from YouTube videos to generate pseudo-labels from a non-streaming teacher model. These labels are then used to train the student streaming model. This technique resulted in an improvement in absolute WER between 1.5% and 9.5%. It is important to note, however, that the languages they use (Spanish, French, Italian and Portuguese) are not considered low-resource. Xu et al. (2022) present two novel techniques of self-distillation where the student model is itself used to leverage its own information better. The first technique is dubbed Neighboring Feature Self-Distillation (NSFD) where for every two neighboring layers, the deeper one is distilled into the shallow one, while the second technique, called Attention-based Feature Self-Distillation (AFSD) introduces an attention mechanism that calculates similarity weights to fuse deep features. These methods achieve a reduction of 6.0% and 8.3% in relative CER, respectively. For Arabic speech recognition, Waheed et al. (2024) investigate pseudo-label distillation methods Gandhi et al. (2023a) across various Modern Standard Arabic (MSA) and dialectal datasets. Their exploration demonstrated the efficacy of distillation in enhancing both efficiency and performance, showcasing that smaller dedicated models can outperform larger multilingual ones. While the pseudo-labeling approach has proven effective across diverse languages Gandhi et al. (2023a); Waheed et al. (2024), it traditionally requires labeled data for filtering low-quality pseudo-labeled examples. In our work we address this limitation by introducing unsupervised data filtering methods, thereby eliminating the dependency on labeled data.
3 Methodology
3.1 Knowledge Distillation
Knowledge distillation is the process through which a small student model learns the behavior of a bigger teacher model Hinton et al. (2015); Sanh et al. (2020); Kim and Rush (2016). Gandhi et al. (2023b) introduce pseudo-label distillation, which generates (English) predictions from a teacher Whisper model and filters them through a WER threshold to only keep the most accurate ones. These pseudo-labels are then used to train a smaller student model. Waheed et al. (2024) build on this approach and show that a Whisper student model beats its teacher model in its average performance in an Arabic multi-dialectal setting. We follow this standard student-teacher framework to distill Whisper into small yet powerful models. The objective of the distillation process can be stated as:
$\mathcal{L}_{KD}=\alpha_{KL}\mathcal{L}_{KL}+\alpha_{PL}\mathcal{L}_{PL}$ (1)
where:
  • $\mathcal{L}_{KL}$ represents the Kullback-Leibler (KL) divergence loss, measuring the difference between the distributions of the teacher and student models. It encourages the student model to mimic the probability distribution of the teacher model.
  • $\mathcal{L}_{PL}$ is the objective to train the student model with pseudo labels as ground truth.
The coefficients $\alpha_{KL}$ (0.8) and $\alpha_{PL}$ (1.0) adjust the relative importance of each loss component in the overall knowledge distillation loss $\mathcal{L}_{KD}$.
3.2 Data Filtering
We employ various methods to filter data without relying on ground truth. These methods are based on the teacher model’s logits, synthetic speech for pseudo-labels, proxy references, and multimodal embeddings for speech and pseudo-labels. We describe each method subsequently.
3.2.1 Proxy Models
This method utilizes a separate, pre-trained ASR model (referred to as proxy model) to generate reference transcripts for the input speech. The quality of the teacher model’s pseudo-label is then assessed based on the Word Error Rate (pWER) between the proxy model’s reference and the teacher’s reference. Lower pWER values indicate better agreement between the teacher and proxy models, suggesting higher confidence and reliability in the pseudo-label. A pre-defined WER threshold can be used to filter out examples exceeding the acceptable error rate. In our work, we use SeamlessM4T-large-v2 as our choice of proxy model.
3.3 Uncertainty Quantification
We attempt to leverage the uncertainty in the teacher’s output to filter out low-quality examples. Measures derived from the teacher model’s confidence scores are used to identify pseudo-labels with high reliability. Two common uncertainty metrics are employed:
Entropy. Entropy measures the information uncertainty associated with the teacher’s predicted probability distribution over the vocabulary. High entropy indicates the model is unsure about the correct word, suggesting a lower confidence in the pseudo-label. The entropy $H$ for a single utterance with $N$ words is calculated as:
$H=-\sum_{i=1}^{N}p_{i}\log_{2}(p_{i})$ (2)
where $p_{i}$ represents the teacher model’s predicted probability for the $i^{\text{th}}$ word in the utterance.
Geometric Mean of Confidence Scores. This metric calculates the geometric mean of individual confidence scores assigned by the teacher model for each word in the pseudo-label. A high geometric mean suggests that the model is confident in all the predicted words, implying a more reliable pseudo-label. The geometric mean $G$ of confidence scores for an $N$-word utterance is:
$G=\sqrt[N]{\prod_{i=1}^{N}c_{i}}$ (3)
where $c_{i}$ denotes the confidence score assigned by the teacher model for the $i^{\text{th}}$ word. During filtering, examples with low entropy and a high geometric mean are considered reliable and are retained for training the student model.
3.3.1 Negative Log Likelihood
We make use of an AceGPT-7B Huang et al. (2024) Language Model to compute the log-likelihood of the teacher’s output. A lower negative log-likelihood indicates higher compatibility between the pseudo-label and the language model’s understanding of natural language, suggesting a more accurate pseudo-label. The negative log-likelihood $NLL$ for a teacher-predicted pseudo-label sequence $y$ of length $T$ is:
$NLL=-\sum_{t=1}^{T}\log(p(y_{t}|y_{1...t-1}))$ (4)
where $p(y_{t}|y_{1...t-1})$ is the probability assigned by the language model to $y_{t}$, the $t^{\text{th}}$ word in the pseudo-label sequence, given the previous words $(y_{1...t-1})$. Pseudo labels with lower negative log-likelihood values are preferred, as they are more likely to form coherent sequences in the language model, indicating higher quality pseudo labels suitable for training the student models.
3.3.2 Multimodal Embeddings
In this approach, SONAR Duquenne et al. (2023) is utilized to generate embeddings for both input speech segments and their corresponding pseudo-labels. These embeddings capture the contextual information of the speech signals and the text representations derived from the pseudo-labels. The similarity between these embeddings is computed using the dot product, serving as a measure of how closely aligned the speech and pseudo-label representations are. A high similarity score indicates strong contextual alignment between the speech and text, suggesting that the pseudo-label accurately reflects the content of the speech segment. This helps in identifying and retaining high-quality pseudo-labels while filtering out those that do not adequately represent the speech content. By leveraging the similarity scores derived from SONAR embeddings, we effectively enhance the quality of the pseudo-labels used in training, improving the overall performance as a result. This method enables more accurate model training in scenarios where labeled data is limited or unavailable, optimizing the efficiency and effectiveness of the learning process.
3.3.3 Similarity with Synthetic Speech
In this approach, we start by generating synthetic speech from the pseudo-labels’ text using XTTS-v2222https://huggingface.co/coqui/XTTS-v2. Next, we compute the similarity between the synthetic speech and the ground truth speech. To assess the quality of the synthetic speech compared to the original speech segment, we utilize metrics such as Perceptual Evaluation of Speech Quality (PESQ) and Mel-Cepstral Distortion (MCD). A high similarity score suggests that the pseudo-labels are likely to have minimal errors. When evaluating these metrics in relation to the Word Error Rate (WER), we observe that PESQ shows a stronger correlation. Therefore, we select PESQ as the primary metric for our experiments. This choice ensures an effective assessment of synthetic speech quality and alignment with the original speech, enhancing the reliability of our pseudo-labeling approach.
3.4 Training Data
We follow Waheed et al. (2024) for the data mixture to train our models. As a result, we randomly select segments from a diverse set of datasets, including MGB2 Ali et al. (2019a), MGB3 Ali et al. (2017), FLEURS Conneau et al. (2022), CommonVoice 15.0 Ardila et al. (2020), QASR Mubarak et al. (2021), the Arabic Speech Corpus Halabi (2016), and the Massive Arabic Speech Corpus (MASC) Al-Fetyani et al. (2021). Specifically, we sample 100K and 500K segments, equivalent to approximately 100 and 500 hours of pseudo-labeled speech data, respectively. Our dataset compilation strictly includes the train splits only from each source. We filter out roughly 27% bad quality examples from our training data using the metrics described in Section 3.
3.5 Teacher and Student Models
In line with previous work Gandhi et al. (2023a); Waheed et al. (2024), we utilize the whisper-large-v2 checkpoint for both pseudo-labeling and as the teacher model during training. We train two variants of students, differing in the number of layers removed from the teacher model. Following (Gandhi et al., 2023b; Waheed et al., 2024), we initialize the student models with maximally spaced layers in the encoder and decoder block of the teacher model. We refer to the models with 16-16 and 32-16 encoder-decoder blocks experimented with 100K segments as UDW-16-16 and UDW-32-16, respectively. The models trained with 500K segments are referred to as UDW-16-16++ and UDW-32-16++.
4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to include the SADA Alharbi et al. (2024) dataset. We describe these corpora below.
4.1 Dataset
FLEURS. The Few-shot Learning Evaluation of Universal Representations of Speech benchmark (FLEURS) Conneau et al. (2022) is an n-way parallel multi-lingual speech dataset and an extension of the FLoRes-101 benchmark Goyal et al. (2022). It comprises 102 languages, each represented by approximately 12 hours of speech. FLEURS is widely utilized for tasks such as ASR, language identification, and speech-to-speech or speech-to-text translation. We use the Arabic split which has MSA speech delivered by speakers with Egyptian accents.
Common Voice. Common Voice (CV) Ardila et al. (2020) is another multilingual dataset powered by volunteer translators and speakers. CV’s latest release to date (CV17.0) counts 124 languages and a total of 31,176 hours. Its Arabic split includes 156 hours of mainly MSA speech.
Multi-Genre Broadcast. Multi-Genre Broadcast (MGB) Ali et al. (2019a, 2017, b) is an Arabic dataset released under three different editions: MGB2 contains 1,200 hours of speech collected from the TV channel Aljazeera Arabic. It is majoritarily MSA speech (> 70%) and is categorized as interviews, reports, and conversations. MGB3, on the other hand, is entirely dialectal with 6 hours of Egyptian speech collected from various YouTube channels. The last edition, MGB5, contains 14 hours of speech extracted from YouTube in the Moroccan dialect.
SADA. The Saudi Audio Dataset for Arabic (SADA) Alharbi et al. (2024) is an Arabic dataset comprised of 668 hours of speech (435 of which are labeled) in both spontaneous and read styles. Most of the speech is in different Saudi dialects (Najdi, Hijazi and Khaleeji among others), but it also includes MSA and dialects from other regions, namely the Levant, Yemen, and Egypt. We report the statistics about the SADA data in Table 1
Dialect Test(S/D) Valid (S/D)
Najdi 1703/2.0709 2249/3.3155
MTOS 1320/4.8044 1048/3.82
Khaliji 1150/1.1308 679/0.6317
Hijazi 809/1.1202 528/0.6423
Unknown 762/0.8325 489/0.4861
NA 167/0.1341 2/0.0004
MSA 157/0.5406 54/0.1682
Egyptian 96/0.0865 45/0.0524
Shamali 18/0.0243 -
Yemeni 7/0.0052 23/0.0349
Levantine - 19/0.0137
Total 6189/10.75 5136/9.17
Table 1: SADA stats. S is the number of segments and D is the duration (in hours).
4.2 Baselines
We conduct evaluations on various multilingual and monolingual speech recognition models using different varieties of Arabic from the datasets mentioned earlier, encompassing both standard and accented MSA, as well as multiple Arabic dialects. We organize these systems into the following categories:
4.2.1 Supervised Models and Commerical Systems.
We evaluate two supervised baselines and a fine-tuned Whisper model for Arabic ASR. The baselines include Wav2Vec2-XLS-RConneau et al. (2020); Babu et al. (2021), trained on CV8.0, and HuBERTHsu et al. (2021), trained on MGB-3 Ali et al. (2017) and a 5.5-hour Egyptian Arabic speech corpus. We fine-tune the third model, whisper-large-v2, on CV11.0 and MGB-2. All models are evaluated using the test and dev splits of MGB-2 and CV11.0, as detailed in Section 4.1.
4.2.2 Zero-Shot Models
Whisper. Whisper Radford et al. (2023) is a versatile multilingual speech model designed for both speech recognition and translation, including Arabic. We assess four Whisper variants: whisper-small (W-S), whisper-medium (W-M), whisper-large-v2 (W-L-v2), and whisper-large-v3 (W-L-v3), using their default decoding parameters with a maximum sequence length of 225 tokens.
SeamlessM4T. SeamlessM4T models excel at generating high-quality transcripts across languages Communication et al. (2023b), but evaluations on them typically focus on English. To fill this gap, we test three of its versions – seamless-m4t-medium (SM4T-M), seamless-m4t-v1-large (SM4T-L-v1), and seamless-m4t-v2-large (SM4T-L-v2) – under a zero-shot setting for Arabic ASR, using the default parameters from the model’s inference pipeline.
4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we take two variants (in terms of size) of distilled models presented in Waheed et al. (2024). Apart from the reported results, we extend the evaluation on SADA’s test and valid splits. We take 16-16 and 32-16 models trained with a WER threshold of 80 on each of 100K and 500K segments. Waheed et al. (2024) find 80 as the optimal threshold which gives the best results. We also take models trained without any filtering and treat those as a lower bound.
4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless noted otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and use WER and CER as the primary metrics for assessing performance.
For distillation, we select roughly 27% of the top segments based on the metrics, to duplicate the amount used in a supervised setup when $\lambda=80\%$. We refrain from conducting hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023b). The distillation process and its key parameters are detailed in Table 5 in Appendix 1.2.
5 Results and Discussion
Model Size CV15.0 MGB2 MGB3 MGB5 Fleurs In-house Data SADA
ALG JOR PAL UAE YEM
Amazon -/- -/- -/- -/- -/- -/- 83.6/70.2 45.5/25.6 52.4/29.0 58.8/40.8 64.7/43.5 -
[…]
HuBERT 0.31 55.2/18.9 49.6/17.3 25.2/9.5 92.4/45.5 34.9/10.9 96.8/44.3 65.2/23.3 73.8/27.9 83.0/36.7 90.5/38.8 75.6/34.5
W-FT 1.5 35.8/21.9 15.3/8.1 48.9/26.9 101.4/62.3 9.8/3.4 115.5/69.6 67.8/37.2 69.6/35.4 105.9/69.1 107.1/64.8 92.3/65.7
MMS-all 1.0 106.4/80.9 39.3/13.4 75.3/34.6 89.7/45.9 23.8/6.3 100.2/78.0 89.8/55.4 99.9/75.1 100.1/78.1 100.2/76.6 77.6/38.0
[…]
SM4T-L-v1 2.3 19.8/7.3 21.8/10.5 44.4/22.6 89.9/52.1 11.1/5.1 87.9/47.8 50.7/18.8 57.5/23.1 61.8/27.4 72.2/32.5 64.9/37.8
SM4T-L-v2 2.3 11.3/3.5 17.3/8.7 36.2/18.6 89.1/53.7 7.6/4.0 92.1/52.0 41.5/14.6 49.5/17.2 55.9/23.3 69.7/30.7 65.0/41.8
[…]
W-M 0.77 29.8/13.2 33.1/18.5 64.3/39.5 127.7/88.3 16.4/5.1 103.7/69.9 50.5/21.1 58.7/24.7 82.5/52.6 86.8/52.0 99.3/74.1
W-L-v2 1.5 19.6/7.8 26.5/15.3 53.0/33.0 99.2/68.9 11.4/3.6 106.4/71.7 42.3/17.0 51.1/22.3 63.8/38.2 77.3/45.5 69.8/51.2
W-L-v3 1.5 15.8/5.2 15.9/7.6 35.7/17.3 79.8/44.6 9.7/3.2 101.9/65.4 43.6/16.3 53.4/22.7 63.4/32.7 76.1/38.9 67.2/45.6
DW-16-16 0.80 22.1/7.2 26.0/10.8 50.5/25.1 82.4/43.3 18.8/6.6 83.0/38.5 50.4/18.2 61.0/23.3 64.6/27.7 72.7/31.6 66.7/38.9
DW-32-16 1.12 18.8/5.9 21.1/8.9 43.8/21.4 78.9/40.4 14.2/4.8 79.5/33.4 44.4/14.7 55.0/19.5 58.1/22.8 68.5/28.1 72.3/49.2
[…]
DW-32-16++ 1.12 17.1/5.5 19.7/8.8 40.7/20.3 76.6/40.6 11.1/3.1 74.6/33.3 41.6/13.4 51.4/18.8 53.5/21.1 63.5/26.8 60.3/35.8
No-filter
$\quad$-- DW-16-16 0.80 22.8/7.6 26.1/11.2 54.1/29.7 95.1/59.1 17.6/6.0 93.4/51.6 51.8/20.2 64.7/27.3 68.0/34.0 78.3/38.8 78.4/49.6
$\quad$-- DW-32-16 1.12 21.2/7.3 22.8/10.4 51.3/30.8 90.5/58.8 14.9/4.9 87.3/63.2 47.6/20.0 57.1/24.9 63.9/35.6 73.0/50.9 70.7/53.6
UDW-16-16 (0.80)
$\quad$-- nnll 23.59/8.15 26.43/11.26 54.37/27.98 92.25/55.25 18.49/6.26 82.44/41.4 63.33/25.7 52.97/20.52 72.03/35.96 93.4/49.2 84.3/55.0
$\quad$-- pesq 24.28/8.41 28.26/12.11 54.16/27.69 93.72/54.88 19.71/6.89 81.54/40.34 64.92/27.41 52.94/20.16 69.67/32.55 86.96/44.17 79.7/50.1
$\quad$-- entropy 23.84/8.1 27.41/12.17 57.29/31.24 94.75/56.64 18.16/6.4 90.05/48 54.76/22.81 64.26/27.67 73.72/37.85 96.35/52.56 92.0/61.8
$\quad$-- conf 23.47/7.83 27.2/11.87 52.66/27.85 87.79/50.73 17.66/6.12 89.15/43.94 53.12/20.29 63.44/25.52 69.03/31.75 79.11/39.27 78.6/49.2
$\quad$-- proxy 22.54/7.48 25.48/11.39 49.41/26.36 84.75/49.97 17.59/7.5 74.53/42.15 61.07/23.69 50.5/19.66 65.43/30.93 84.12/41.94 68.6/46.2
$\quad$-- sonar 24.06/8.04 28.34/11.86 55.09/28.66 85.59/49.21 20.42/7.06 86.93/43.48 56.36/22.61 65.4/27.43 70.7/32.89 76.48/36.13 74.9/45.6
[…]
$\quad$-- nnll 18.92/6.24 22.63/10.12 48.08/25.39 94.07/55.53 13.29/4.47 75.43/35.85 55.24/20.88 44.69/16.04 63.71/30.49 83.81/41.38 73.4/46.1
$\quad$-- pesq 21.7/7.5 24.2/10.6 49.0/26.4 87.0/51.0 16.1/5.3 88.8/43.2 47.5/17.2 58.5/22.7 67.0/30.9 75.5/36.4 73.3/48.1
$\quad$-- entropy 19.09/6.53 21.82/10.34 46.93/28.71 87.11/66.87 13.29/4.34 90.99/84.02 46.68/21.07 58.16/31.08 64.94/44.23 76.35/54.25 68.3/52.3
$\quad$-- conf 20.27/6.46 22.76/9.79 46.6/23.42 83.43/48.61 14.53/4.73 73.42/36.08 57.25/21.99 61.06/25.82 47.5/17.7 85.71/42.82 66.2/41.3
$\quad$-- proxy 19.04/6.17 21.51/9.87 44.29/22.45 82.64/46.05 14.23/6.23 80.36/36.11 45.46/15.6 56.01/20.88 61.63/25.69 69.52/29.49 64.6/41.9
$\quad$-- sonar 17.82/5.62 21.1/8.98 45.29/23.4 80.39/46.97 13.07/4.41 78.97/37 44.3/15.11 54.34/19.56 58.76/24.29 66.82/28.24 58.7/35.6
[…]
$\quad$-- proxy 15.5/4.8 23.3/10.4 46.8/24.3 84.6/48.0 14.2/4.8 82.9/41.6 46.4/16.3 57.0/21.1 61.1/27.5 70.8/33.0 60.2/35.3
$\quad$-- sonar 18.9/6.1 22.8/9.8 47.2/24.2 82.0/47.1 14.9/5.0 81.5/38.9 47.7/17.2 57.4/21.4 62.1/26.2 69.2/29.5 59.4/33.96
UDW-32-16-++ (1.12)
$\quad$-- proxy 18.1/5.8 20.8/9.2 43.1/21.1 82.3/44.1 12.6/4.2 80.8/37.1 43.1/14.5 54.4/20.1 57.7/23.1 68.5/29.2 55.1/31.4
$\quad$-- sonar 17.0/5.3 21.9/9.9 44.1/22.4 79.8/44.2 12.8/3.9 78.2/34.6 54.2/19.2 44.4/15.1 58.2/23.2 66.6/27.0 61.6/33.4
Table 2: WER/CER scores after normalization and removing diacritics. All baseline distilled models (dw-) are trained with a filtering threshold of 80 if not specified. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, U - Unsupervised, D - Distil.
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. We do the distillation with two variants of student models. The first student has 16 of each of the encoder and decoder blocks (UDW-16-16) and the second one has 32 encoder and 16 decoder blocks (UDW-32-16). Following Waheed et al. (2024), we experiment with 100K and 500K segments. We report our main results in Table 2, and the averages across different evaluation setups in Table 3. Results on the top five varieties [^3] We refer dialect labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 4. The reported results are computed on normalized text that includes removing diacritics. Results on the orthographic text (no normalization) can be found in Appendix 1.2.
Waheed et al. (2024) show that their supervised models perform well on in-distribution evaluation, however, their performance degrades considerably under novel and challenging conditions. Our evaluation of these supervised models shows similar results. We find HuBERT trained on MGB3 Ali et al. (2017) corpus outperforms all other supervised baselines across different evaluation setups.
In zero-shot models, we compare our results with MMS Pratap et al. (2023a), three variants of SeamlessM4T family models (medium, large-v1, large-v2), and four Whisper family models. We extend the evaluation to SADA data for these models as well. While larger variants of both SeamlessM4T and Whisper perform on standard benchmark setups like Common Voice and Fleurs they perform poorly on the novel and challenging settings. For example, the best zero-shot model SeamlessM4T-large-v2 gives just 7.6% WER on Fleurs compared to 65.0% WER on the SADA test set as shown in Table 2. This underscores the inadequate evaluation and limited generalization capability of zero-shot models in unseen and challenging settings.
Further, we take the reported results of distilled models in Waheed et al. (2024) and extend its evaluation to SADA data. As demonstrated in the reported results, the best-distilled model DW-32-16++ outperforms all other models including its teacher (large-v2), best of Whisper (large-v3), and best of baselines (SeamlessM4T-large-v2).
We trained two versions of teacher models using datasets of 100K and 500K segments, respectively. Approximately 27% of the examples were removed from each dataset based on specific metrics, aligning our data set up with the one used in Waheed et al. (2024), which employs a filtering threshold of 80. Consequently, we consider models trained with this threshold of 80 to represent an upper bound, while models trained on the original, unfiltered pseudo labels serve as our baseline for comparison.
From our experiments, we observed that two of our best measures, sonar-sim and proxy-ref based filtering, perform comparably to or even surpass the equivalent supervised setups. For instance, the model DW-32-16 trained on 100K examples with a filtering threshold of 80 achieved a WER of 35.3% on the benchmark test split, 66.6% on the SADA test split, and 61.1% averaged om five novel dialects. By comparison, our best-performing model in a similar configuration, UDW-32-16-sonar, achieves 35.5% WER on the benchmark test split, 58.7% on the SADA test, and 60.6% on five dialects. This reflects an almost 3% improvement in average score compared to the supervised setup (51.6 vs 54.3) and more than a 7% improvement over the no-filter setup, as detailed in Table 3.
Additionally, our experiments show that the smaller student model (UDW-16-16) performs better with the proxy-ref as a filtering measure. More specifically, the UDW-16-16, experimented with 100K examples and utilizing proxy-ref to filter out low-quality pseudo labels, matches the results of supervised setups (58.5% vs 57.6%) and shows over a 6% improvement compared to setups where pseudo labels are not filtered (64.2%).
Model Bench SADA2022 IH
Test Valid Test Valid
HuBERT 51.4/20.4 54.4/22.7 75.6/34.5 73.7/31.9 81.9/34.2
W-FT 42.2/24.5 47.4/28.6 92.3/65.7 81.7/56.2 93.2/55.2
[…]
SM4T-v2 32.3/17.7 34.7/19.4 65.0/41.8 63.2/40.8 61.7/27.6
W-M 54.3/32.9 59.6/38.0 99.3/74.1 91.9/66.7 76.4/44.1
W-L-v2 41.9/25.7 46.8/29.7 69.8/51.2 66.5/48.5 68.2/38.9
W-L-v3 31.4/15.6 33.0/17.1 67.2/45.6 58.4/39.3 67.7/35.2
DW-16-16 40.0/18.6 42.2/20.5 66.7/38.9 65.2/37.8 66.3/27.8
DW-32-16 35.3/16.3 37.7/17.9 66.9/47.3 64.5/43.9 61.1/23.7
DW-16-16++ 36.7/17.8 38.8/19.4 69.9/40.6 64.7/36.6 62.3/27.2
DW-32-16++ 33.0/15.7 34.9/17.1 60.3/35.8 57.1/33.4 56.9/22.7
[…]
$\quad$-- DW-16-16 43.1/22.7 46.5/25.1 78.4/49.6 72.5/45.9 71.2/34.4
$\quad$-- DW-32-16 40.1/22.5 43.6/25.8 70.7/53.6 66.0/45.0 65.8/38.9
UDW-16-16
$\quad$-- proxy 40.0/20.5 42.3/22.1 68.6/46.2 65.4/42.2 67.1/31.7
$\quad$-- sonar 42.7/21.0 44.9/22.8 74.9/45.6 71.2/43.5 71.2/32.5
[…]
$\quad$-- proxy 36.3/18.2 39.1/19.6 64.6/41.9 58.7/36.1 62.6/25.6
$\quad$-- sonar 35.5/17.9 37.8/19.7 58.7/35.6 56.4/34.7 60.6/24.8
[…]
$\quad$-- proxy 36.9/18.4 41.4/21.2 64.1/39.2 60.2/35.3 63.6/27.9
$\quad$-- sonar 37.2/18.4 39.3/19.9 61.7/35.4 59.4/34.0 63.6/26.6
[…]
$\quad$-- proxy 35.1/17.1 37.2/18.7 58.9/33.6 55.1/31.4 60.3/23.8
$\quad$-- sonar 35.4/16.9 37.7/18.5 57.9/33.1 54.7/31.3 60.9/24.8
Table 3: Average WER/CER across different evaluation datasets. WER/CER scores after normalization and removing diacritics.
Scaling the Data.
Following Waheed et al. (2024), we also expand our dataset from 100K to 500K segments, equivalent to approximately 500 hours of data. As with our other experiments, we filter out approximately 27% of low-quality examples from our training data to align with the quantity used in supervised baselines. We train two models, UDW-16-16 and UDW-32-16, employing proxy-ref and sonar-sim as filtering criteria.
As shown in Table 3, our UDW-32-16++ model, trained with 500K segments, performs comparably with the supervised counterpart DW-32-16++. For instance, UDW-32-16++ with proxy reference filter achieves WER of 51.4%, whereas DW-32-16++ achieves 50.0% WER. Notably, our smaller student model on 500K segments (UDW-16-16++) results in superior performance compared to models trained under a supervised filtering setup. For example, UDW-16-16++ with sonar-sim gives 53.9% WER compared to 56.2% WER for DW-16-16++. These results showcase the efficacy of our data filtering methods.
Generalization to Unseen Dialects.
Model Split NJD MTOS KHLJ HJZ UNK
W-FT Test 106.4/77.5 77.0/51.8 117.8/85.4 84.2/61.5 139.9/112.2
Valid 77.1/52.6 63.4/41.1 139.4/100.3 119.1/89.7 140.3/107.6
SM4T-v1 Test 56.5/30.9 74.0/46.0 61.9/32.2 54.6/29.0 69.6/39.4
Valid 51.9/28.1 68.7/44.0 61.7/31.5 54.2/30.9 62.3/35.2
SM4T-v2 Test 52.6/31.1 76.2/53.1 58.2/30.4 53.3/32.0 74.7/45.1
Valid 52.2/30.7 75.8/53.7 65.1/35.3 51.1/30.3 59.8/34.4
W-M Test 88.2/65.8 107.9/79.3 102.8/77.0 86.7/59.7 134.0/122.2
Valid 80.4/56.9 102.8/75.1 89.5/62.9 72.9/52.0 127.7/106.5
W-L-v2 Test 56.6/39.9 74.0/57.4 80.2/54.4 63.4/39.6 98.2/80.7
Valid 60.9/41.4 72.9/55.4 67.7/44.9 64.5/43.6 68.0/67.1
W-L-v3 Test 53.1/31.6 73.7/53.7 70.5/44.1 61.7/38.6 89.0/61.3
Valid 49.3/30.2 65.5/47.7 67.5/39.2 46.5/27.2 67.7/49.2
DW-16-16 Test 58.2/30.8 74.9/47.6 65.1/32.7 58.2/30.7 68.4/39.8
Valid 59.4/31.4 70.6/44.7 66.2/35.2 61.1/32.8 69.9/39.8
DW-32-16 Test 58.7/35.8 74.9/60.1 66.7/38.7 58.7/34.1 70.1/44.5
Valid 58.3/34.8 69.7/54.1 67.5/37.6 62.7/38.2 68.3/40.1
DW-16-16++ Test 58.8/30.9 81.9/50.7 62.1/31.5 57.5/31.0 77.5/46.8
Valid 56.8/29.8 72.0/43.8 62.3/31.8 60.2/33.0 75.2/41.0
DW-32-16++ Test 51.7/28.3 67.8/43.1 57.5/29.4 52.6/28.6 67.5/41.3
Valid 50.3/27.3 61.8/38.3 62.3/34.5 53.7/28.1 66.4/43.0
[…]
$\quad$-- DW-16-16 Test 61.6/34.8 90.4/59.7 72.9/41.4 68.5/42.3 88.8/63.0
Valid 64.8/38.9 80.3/53.4 71.4/41.9 65.5/37.7 77.0/54.3
$\quad$-- DW-32-16 Test 58.2/42.8 82.3/63.9 67.3/47.0 60.4/45.9 76.6/63.2
Valid 57.9/35.2 73.7/54.9 68.6/43.3 56.8/36.5 72.3/49.6
UDW-16-16
$\quad$-- proxy Test 60.2/35.5 75.6/55.6 65.7/38.9 62.1/39.6 75.6/52.0
Valid 59.3/34.0 70.7/50.9 66.5/39.1 61.4/37.1 68.7/41.2
$\quad$-- sonar Test 64.6/35.8 59.9/30.3 84.6/55.7 72.1/38.8 66.6/36.8
Valid 64.8/35.9 67.5/39.3 78.1/52.4 69.9/38.7 65.3/36.7
[…]
$\quad$-- proxy Test 53.4/31.1 75.7/54.0 60.0/32.1 51.4/30.8 72.8/46.0
Valid 51.0/29.3 65.9/44.3 58.1/29.1 53.7/28.6 64.9/36.6
$\quad$-- sonar Test 48.8/25.4 52.6/23.6 67.5/45.9 58.3/29.9 48.9/25.5
Valid 49.2/26.0 51.6/26.7 62.5/44.1 58.9/30.3 52.6/29.5
[…]
$\quad$-- proxy Test 53.4/29.7 73.9/48.8 61.9/33.8 53.3/29.6 68.6/42.8
Valid 52.7/27.8 66.6/42.0 62.0/34.3 55.7/32.2 67.3/41.7
$\quad$-- sonar Test 53.7/28.4 69.4/43.3 61.4/30.8 53.7/27.5 64.6/37.0
Valid 53.6/27.5 64.4/40.3 62.3/32.4 55.7/30.7 63.7/35.8
UDW-32-16++
[…]
Valid 49.0/25.3 60.4/37.4 57.6/30.1 50.7/25.6 60.7/37.4
$\quad$-- sonar Test 50.5/26.3 65.3/40.8 57.3/28.3 49.1/24.8 61.8/34.5
Valid 49.3/25.3 58.8/37.0 58.1/30.2 53.6/29.8 61.1/34.2
Table 4: Results on top five dialects/categories in SADA data. WER/CER scores after normalization and removing diacritics.
We evaluate our models under various unseen conditions, including five novel dialects that our models have not encountered before. Additionally, we compute category-wise error rates for both the SADA test and valid splits. We present results on top five categories  [^4] We refer fategories as dialectal and non-dialectal labels in SADA data. As detailed in Table 4, our best model consistently outperforms all others when averaged across these novel dialects.
Furthermore, our top-performing model, UDW-32-16++, demonstrates superior performance over DW-32-16++ in the top five SADA categories. For instance, using proxy reference as a filtering measure, UDW-32-16++ achieves 58.06%, compared to DW-32-16++’s 59.42% WER averaged across top five categories in SADA test split. This highlights our ability to distill smaller models from large Whisper models, maintaining or improving performance while reducing size, all without requiring labeled data.
Effectiveness of Unsupervised Metrics to Filter Low-Quality Pesudo Labels.
We investigate the effectiveness of several of these metrics for filtering low-quality pseudo-labels in speech recognition tasks, specifically targeting instances with a WER higher than 80%, 40%, and 20%. To assess their efficacy, we calculate the area under the curve (AUC) as we show in Figure LABEL:fig:side-by-side for detecting low-quality examples. The results indicate that sonar-sim achieves an AUC of 0.77 for detecting examples with $WER>80$, demonstrating reasonably high discriminative power in identifying low-quality labels. The proxy-ref metric shows slightly better performance, with an AUC of 0.82, indicating robust capability in distinguishing between high and low-quality pseudo-labels. In contrast, the confidence-based measure yielded an AUC of 0.68, which shows relatively less discriminative power in filtering out undesirable labels. These findings highlight sonar embeddings and the proxy reference-based measure as promising tools for improving the quality of pseudo-labels in scenarios where ground truth data is unavailable.
6 Conclusion
In this study, we explore methods for distilling large Whisper models into smaller, more efficient ones. Our proposed techniques for filtering out low-quality examples eliminate the need for labeled data, bridging a gap in prior research and facilitating the creation of compact and effective speech recognition models for settings with limited labeled data. We conducted a comprehensive evaluation of our trained models, comparing them against various strong zero-shot and distilled models. Our results demonstrate that our models can outperform both the teacher model and models distilled with supervised data filtering setups.
We evaluate our models across a diverse range of Arabic varieties and dialects, assessing their generalization and adaptability to linguistic diversity. From large multilingual speech models (Whisper), we distill language-specific models optimized for Arabic ASR. Our evaluation spans three different setups, encompassing a wide array of linguistic varieties (ie: dialects), and reveals that despite their reduced size (25-50%), these distilled models consistently outperform their base counterparts and can compete with state-of-the-art models twice their size—all achieved without the need for labeled data.
Notably, our model-based unsupervised data filtering methods (proxy and sonar) demonstrate superior robustness across linguistic variations. Moving forward, we aim to explore model-free approaches to further enhance the efficacy of filtering low-quality pseudo labels. Additionally, our future research will extend to include extremely low-resource languages and domains.
7 Limitations
In this study, we distill small Whisper models from relatively large ones via pseudo-labeling and unsupervised data filtering. Our distilled models are computationally efficient and maintain a performance similar to or better than the base teacher model and models trained in a supervised data filtering setup. Unlike Waheed et al. (2024); Gandhi et al. (2023a), our approach does not utilize any labeled data in the distillation process, making it directly applicable in data-scarce settings. However, despite these advantages, we acknowledge several limitations in our work, which we outline below.
Efficiency. Our distilled models achieve 25-50% compute efficiency relative to their larger counterparts while maintaining comparable performance. However, the training of these models requires significant computational resources.
Our main approach relies heavily on a robust reference model to serve as a proxy for filtering lower-quality pseudo labels. Specifically, we utilize SeamlessM4T-large-v2, a state-of-the-art model with 2.3 billion parameters, to generate proxy references which is then used to filter out low-quality data points. For similarity-based measures, we use SONAR to generate multimodal embeddings from speech and pseudo labels. These embeddings provide contextual similarity which is then utilized to discard low-quality pseudo labels. We use AceGPT (7B), to compute the log-likelihood of the pseudo labels is which leveraged to filter out low-quality examples.
Although with these measures we attain performance on par or better than the supervised setup, it’s important to highlight that each of these methodologies entails additional computational overhead.
Multilinguality. We use SeamlessM4T-large-v2 for generating proxy references, SONAR for generating multimodal embeddings, AceGPT (7B) for computing log-likelihood, and XTTS-v2 for generating synthetic speech. The multilingual capabilities of these models are crucial for effectively applying our techniques to a wide range of languages and dialects. However, a significant limitation of our approach is that it is constrained to languages supported by these models. This dependency restricts our ability to extend our distillation process to languages beyond the scope of the models’ multilingual capacities.
Evaluation. Arabic is a linguistically rich and complex language with over 400 million speakers Abdul-Mageed et al. (2021), resulting in its wide range of varieties and dialects. We evaluate all the models on ten different datasets representing different varieties, including five novel dialects collected and curated by native speakers and never seen before by any models. However, our varieties do not cover all Arabic-speaking regions. We aim to address this in future work by covering more varieties and dialects.
Distillation Training Data. We distilled four variants of student models using 100K and 500K segments of which approximately 25% are filtered. We see improvement going from 100K ($\approx$100100100100 hours) to 500K ($\approx$500500500500 hours) segments. As Gandhi et al. (2023b) shows going over 1000 hours results in a better model, we aim to study how distillation can be done under a low resource setting which is why we do not scale the data. Additionally, we also keep the WER threshold high (80) so that we remain close to a setting where no labeled data is available (even for filtering). It would be interesting, however, to see how distilled models may perform on unfiltered data in low-resource setting.
Nature of Speech Data. Despite putting together a never-seen dataset of under-represented Arabic dialects, we realize that sourcing our data from television series renders its nature distant from speech spoken in the wild. This type of content tends to be more “theatrical” and involves different elements such as background music and laughing tracks that do not accurately reflect regular conversational Arabic. Consequently, this could fail to accurately portray the performance of these models on real speech.
Appendix A Appendix
1.1 Training Parameters
Parameter Value
warmup_steps $50$
learning_rate $0.0001$
lr_scheduler_type constant_with_warmup
batch_size $128$
max_label_length $225$
gradient_accumulation_steps $1$
dtype bfloat16
Table 5: Training parameters. All the training parameters are the default ones provided in Huggingface Seq2SeqTrainingArguments unless specified otherwise in this table.
1.2 Results
Model Size CV15.0 MGB2 MGB3 MGB5 Fleurs In-house Data SADA
ALG JOR PAL UAE YEM
Amazon -/- -/- -/- -/- -/- -/- 88.0/71.6 59.2/29.1 63.4/32.2 71.1/44.3 $7$
[…]
HuBERT 0.31 76.5/31.0 59.4/20.3 43.3/16.5 95.0/48.7 48.9/14.4 96.2/45.6 70.6/25.4 81.5/31.4 87.9/39.9 91.3/40.8 81.3/37.1
W-FT 1.5 70.0/33.8 29.4/10.9 60.1/32.2 105.0/64.3 28.7/7.3 114.5/70.3 75.1/39.0 81.3/38.7 113.7/70.9 110.1/65.6 101.4/67.6
[…]
SM4T-L-v1 2.3 44.2/19.1 25.9/11.7 52.5/27.6 92.8/55.9 22.6/7.6 89.7/50.3 59.1/21.7 64.7/25.8 69.0/30.3 81.5/37.0 72.4/40.8
SM4T-L-v2 2.3 37.7/15.8 22.4/9.9 46.7/23.9 92.1/58.4 19.8/6.5 94.8/55.2 51.3/17.6 58.5/20.1 65.6/26.9 80.6/35.5 72.2/44.4
[…]
W-M 0.77 55.1/24.2 37.6/19.6 71.5/43.7 129.7/89.4 24.0/7.1 103.9/71.4 59.0/23.9 66.8/27.6 90.7/55.7 95.2/56.2 106.0/76.3
W-L-v2 1.5 46.9/19.6 33.7/16.9 60.6/37.7 101.1/71.1 19.7/5.6 106.9/74.6 51.2/19.6 60.2/25.2 73.2/41.2 86.9/50.1 78.0/53.5
W-L-v3 1.5 43.2/16.9 20.4/8.6 44.6/22.5 82.0/47.7 16.4/4.8 103.8/68.9 52.7/18.9 64.3/26.4 72.3/35.9 86.0/43.3 74.6/47.9
DW-16-16 0.80 48.0/18.9 33.2/12.5 57.1/29.6 84.1/46.2 26.2/8.5 83.8/40.2 57.8/20.5 68.2/26.2 72.0/31.0 80.0/35.6 72.0/40.9
DW-32-16 1.12 45.6/17.7 27.7/10.3 51.2/26.1 80.9/43.4 22.0/6.6 80.5/35.1 52.6/17.1 62.9/22.4 66.7/26.3 77.3/32.6 66.9/47.3
[…]
DW-32-16++ 1.12 44.7/17.3 25.2/10.0 48.8/25.2 79.0/43.7 20.2/5.0 76.4/35.4 50.0/15.9 60.1/21.8 63.2/24.7 73.5/31.5 67.2/38.1
[…]
$\quad$-- DW-16-16 0.0 48.3/19.1 34.2/13.0 60.2/33.9 96.8/61.3 24.9/7.9 93.8/53.1 58.9/22.4 72.3/30.2 75.5/37.2 84.8/42.3 83.9/51.6
$\quad$-- DW-32-16 0.0 47.2/18.8 29.0/11.8 58.3/35.0 92.5/60.8 23.5/7.0 88.0/64.2 55.3/22.4 65.5/27.9 72.5/38.8 80.4/53.8 76.7/55.5
UDW-16-16
$\quad$-- nnll 0.0 49.12/19.55 32.18/12.64 60.76/32.4 94.05/57.88 25.75/8.09 88.88/44.93 70.14/28.53 59.84/22.8 78.73/38.95 93.3/50.73 89.2/56.9
$\quad$-- pesq 0.0 49.46/19.73 34.27/13.55 60.58/31.99 95.63/57.22 26.8/8.64 87.91/43.83 71.92/30.19 60.01/22.47 76.7/35.72 87.51/45.81 84.7/52.1
$\quad$-- entropy 0.0 49.03/19.5 33.4/13.57 62.74/35.19 96.45/58.9 25.81/8.29 90.31/49.75 61.72/25.08 71.29/30.42 80.69/40.86 101.99/55.25 96.8/63.7
$\quad$-- conf 0.0 48.73/19.3 34.01/13.44 59.2/32.17 89.77/53.36 24.68/7.86 89.61/45.69 60.05/22.57 71.27/28.52 76.71/35.08 85.44/42.53 84.0/51.3
$\quad$-- proxy 0.0 47.9/18.76 30.96/12.7 56.03/30.8 86.44/52.66 24.69/9.25 81.7/45.74 68.43/26.62 58.08/22.07 72.96/34.2 84.6/43.64 74.1/48.2
$\quad$-- sonar 0.0 48.82/19.4 33.88/13.19 61/33.03 87.08/52.03 27.75/8.91 87.6/45.09 62.72/24.81 71.92/30.21 76.9/35.99 82.49/39.72 78.9/47.3
[…]
$\quad$-- nnll 0.0 45.55/17.88 29.02/11.52 54.66/29.79 96.01/57.91 20.19/6.13 84.03/39.91 63.45/23.8 52.63/18.43 71.98/33.71 84.57/43.01 79.6/48.0
$\quad$-- pesq 0.0 47.5/19.0 32.0/12.4 55.6/30.8 88.9/53.8 23.5/7.1 89.3/44.9 55.2/19.6 66.4/25.6 75.0/34.2 83.3/40.4 79.0/50.0
$\quad$-- entropy 0.0 45.68/18.18 29.86/12.07 53.88/32.97 88.92/68.55 21.55/6.32 91.78/84.74 54.32/23.38 66.47/33.82 73.62/47.04 83.91/56.95 74.7/54.1
$\quad$-- conf 0.0 47.07/18.29 31.12/11.62 54.04/28.14 85.53/51.41 22.78/6.68 81.01/39.84 65.69/25 69.83/29.36 55.32/20.11 86.59/44.52 72.9/43.6
$\quad$-- proxy 0.0 45.38/17.71 28.22/11.32 51.47/27.06 84.54/48.86 21.41/7.93 81.25/37.94 53.42/18.05 63.9/23.76 70.19/29.16 78.38/34.03 71.0/44.1
$\quad$-- sonar 0.0 44.74/17.38 27.35/10.35 52.72/28.08 82.33/49.95 21.05/6.29 80.58/38.85 52.29/17.53 63.17/22.67 67.71/27.87 75.42/32.57 65.0/37.7
[…]
$\quad$-- proxy 0.0 52.0/22.5 28.4/11.6 53.8/28.8 86.5/50.6 21.9/6.7 83.6/43.2 54.2/18.8 64.9/24.1 69.3/30.9 78.9/37.2 66.2/37.3
$\quad$-- sonar 0.0 45.5/17.7 29.9/11.4 53.9/28.8 83.7/50.1 22.8/7.0 82.6/40.6 55.4/19.6 65.0/24.4 70.2/29.6 77.5/33.9 65.7/36.0
UDW-32-16-++
$\quad$-- proxy 0.0 44.9/17.5 26.6/10.5 50.4/25.8 84.4/47.1 20.7/6.1 81.6/39.0 51.5/17.0 62.6/23.0 66.5/26.6 77.6/33.8 62.1/33.6
$\quad$-- sonar 0.0 44.2/17.1 29.0/11.5 51.2/27.0 81.7/47.2 20.6/5.7 79.5/36.5 62.3/22.2 52.5/17.5 67.1/26.8 76.0/31.8 54.7/31.3
Table 6: WER/CER scores on orthographic transcription. Average is the mean score across all the evaluation sets. All distilled models are trained with a filtering threshold of 80. We report the score on the test split of each dataset.
Appendix B Efficacy
This is an appendix.